LABORATORY DETECTION OF THE NEW CARBON CHAIN RADICAL H_{2}CCCCN

\underline{WEI CHEN}, MICHAEL C. McCARTHY, MICHAEL J. TRAVERS, PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138 and Division of Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138; ELAINE W. GOT-TLIEB, CARL A. GOTTLIEB, Division of Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138; MICHAELEEN R. MUNROW, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The cyanomethyl radical, H_{2}CCN, was first found in the interstellar medium and then detected in the laboratory in 1988, however, the longer members of this sequence have remained undetected. We report the first detection of the cyanopropynyl radical, H_{2}CCCCN, in a pulsed-discharge nozzle Fourier transform microwave spectrometer. Four rotational transitions (N=2-1 through 5-4) of both H_{2}CCCCN and H_{2}CCCC^{15}N were measured and the complex fine and hyperfine structure was assigned. These results will be presented and the possibility of detecting the next member of the series will be discussed.